OpenStudy (anonymous):

would you know where these 2 linear equations intersect? if it does? 2x-3y=6 and 6x=18+9y

OpenStudy (anonymous):

they would intesect unless the lines are parellel

OpenStudy (anonymous):

1)2x-3y=6 2)6x-9y=18 =3(2x-3y=6) 2) the second equation is factor of first one so they are parellel

OpenStudy (anonymous):

these are parallel = would not intersect

OpenStudy (anonymous):

how did u know these are parallel lines? can you please show me solutions?

OpenStudy (anonymous):

u there?